OpenStudy (anonymous):

How do you find the value of x and y for 4x+5y=102 7x+5y=126

OpenStudy (anonymous):

Elimination, substitution, matrix, graph? which method?

OpenStudy (anonymous):

Solve the first equation for y, then plug that into the second equation to solve for x. Once you know x you can plug that into the first equation to find y.

OpenStudy (anonymous):

i was thinking substitutio imranmeah

OpenStudy (anonymous):

but how do i solve for y?

OpenStudy (anonymous):

4x+5y=102 7x+5y=126 We will solve by elimination

OpenStudy (anonymous):

We want y to cancel to we multiply one equation by -1 4x+5y=102 -7x-5y=-126 ---------------- -3x=-24 x=8

OpenStudy (anonymous):

Now plug that x value back in to find y value 4(8)+5y=102 32+5y=102 5y=102-32 5y=72 y=14.4

OpenStudy (anonymous):

alright!

OpenStudy (anonymous):

that really helps!

OpenStudy (anonymous):

how does that cancelling work?

OpenStudy (anonymous):

5y and -5y cancel when you add them

OpenStudy (anonymous):

If I helped you may award me a medal or become my fan. Click Good Answer to give me a medal.

OpenStudy (anonymous):

where is good answer?

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

i figured it out XD i had to refresh